Homework 4 (Chapter 4 and 5) 6 Refer to the demand schedule below: 60 1e0 150 200 250 108 400 a. Suppose the price increases from $10 to $20. Demand is Click to selec)and total revenue (Cick to selecty b. Suppose the price increases from $30 to $40. Demand is (Cick to select) and total revenue Click to selec) es C. Suppose the price increases from $50 to $60. Demand is (Click to selectand total revenue (Click to selec) (Click to unit-elastic inelastic K Prev 15 of 23 Next
Answer #1

A) P1= 10, P2=20, Q1=350, Q2=300

Ped= Q2-Q1/Q1/P2-P1/P1

=-50/350/10/10= -1/7 <1

HENCE. PED is inelastic.

2)p1= 30, p2=40, q1=250, q2=200

Ped= -50/250/10/30= -3/5<1

InElastic

C) p1=50, p2=60, q1=150, q2=100

Ped= -50/150/10/50= -25/15=-5/3>1

Elastic

WHEN the MAGNITUDE OF PED IS GRETAER THAN 1 IT IS L

ELASTIC, IRRESPECTIVE OF ITS SIGN. AND WHEN IT IS LESS 1 IT IS INELASTIC.

*drop down for total revenue is not provided in the image*